ST-1 Jldividual costs of capital) Compute the cost for the following sources o (a) A $100 par-value bond with a market price of $97 Costs and a coupon interest rate of 10% e bonds mature in 10 years and the corpo (b) Preference shares selling for $10 with an annual unfranked dividend payment of 80 cents. (c) Internally generated equity totalling $4.8 million. The price of ordinary shares is $7.50 per for a new issue would be approximately 5%. T rate tax rate is 30%. Estimate the cost of debt before tax. If the company sells a new issue, the costs will be 90 cents per share. share, and the dividends per share were 98 cents last year. These dividends are not expected to increase, or to be franked. (d) New ordinary shares whose most recent dividend was 28 cents. The companys dividends per share should continue to increase at an 8% growth rate into the indefinite future. The market price of the shares is currently $5.30; however, issue costs of 60 cents per share are expected if the new shares are issued. Estimate the after-tax cost of ordinary shares, assum- ing that dividends are not franked. e) Preference shares (see b) and ordinary shares (see d) for taxation category I companies under the imputation system. The companys marginal tax rate is 30%.
Answer #1

a)let us find the ytm of bond we use rate formuale in excel
=rate(nper,pmt,pv,fv,type)
nper=10
pmt=100*10%=10
pv=97*(1-5%)=-92.15
fv=100
=rate(10,10,-92.15,100,0,1)
=11.35%
Cost of debt=11.35%*(1-30%)=7.95%

B)Cost of preferred stock= preference dividend/parvalue*(1-flotation costs)
=0.8/(10-0.9)
=8.79%

c)Cost of internal equity= dividend/price +growth
=(0.98/7.5)+0
=13.07%

d)Price*(1-flotation)=Dividend*(1+g)/(k-g)
=(5.3-0.6)=0.28*(1+8%)/(k-8%)
k=14.43%